The length of a rectangular garden is 10 feet longer than its width. If the garden's perimeter is 204 feet, what is the area of the garden in square feet?
Answer by ikleyn(52781)

.

W + (W+10) = 204/2 = 102   (half of the perimeter).


Hence,


2W = 102-10 = 92  ----------->  W = 92/2 = 46.


W = 46;  L = 46+10 = 56.


The area = 46*56 = use your calculator.